A bar of magnetic material consists of atoms with magnetic dipole moments of magnitude μatom = 3.50 ×10-23J/T. The bar contains N = 2.4 × 1021 atoms and has a volume of V = 3.0 × 10-8 m 3 . What is the saturation magnetization? If only 0.300% of the atomic dipoles are aligned, what is the magnetization?
Answer #1

N= 2.4× 10^21

m= 3.5×10^(-23) J/T

V= 3.0×10^(-8) m^3

Effective no. Of atoms

N' = .003× 2.4× 10^(21)

N'= 7.2×10^(-18)

Total effective dipole moment

m'= N'×m/V

= 7.2×10^(18)×3.5×10^(-23)/(3.0×10^(-8))

= 8.4× 10^(3) J/(T-m^(3))